OpenStudy (anonymous):

Jerry invested $7000, part at 8% simple interest and the rest at 5% simple interest for a period of 1 year. If he received a total annual interest of $476 from both investments, how much did he invest at each rate?

OpenStudy (anonymous):

0.08x+0.05(7000-x)=476 0.08x+350-0.05x=476 0.03x=126 x=4200 at 8% 7000-x=2800 at 5%